OpenStudy (anonymous):

solve for x over the complex numbers. write in a+bi form. 52=4-2(x-3)(x-3)

OpenStudy (anonymous):

Well, for starters I would try and get x by itself. And then find what x=. Then I would concentrate on the second part which is putting x into a+bi form.

OpenStudy (anonymous):

i can get it down to x=3+- 24i . but im stuck on simplifying it any further

OpenStudy (campbell_st):

rewrite the equation and start by subtraction 4 from both sides 48 = -2(x-3)^2 that should make it a little easier...

OpenStudy (campbell_st):

rewrite the radical \[\sqrt{-24} = 2i \sqrt{6}\]

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

it did thanks